# POPGym Arcade: Parallel Pixelated POMDPs

### v1

@@ Abstract
We introduce POPGym Arcade, a benchmark consisting of 7 pixel-based environments each with three difficulties, utilizing a single observation and action space. Each environment offers both fully observable and partially observable variants, enabling counterfactual studies on partial observability. POPGym Arcade utilizes JIT compilation on hardware accelerators to achieve substantial speedups over CPU-bound environments. Moreover, this enables Podracer-style architectures to further increase hardware utilization and training speed. We evaluate memory models on our environments using a Podracer variant of Q learning, and examine the results. Finally, we generate memory saliency maps, uncovering how memories propagate through policies. Our library is available at https://github.com/bolt-research/popgym_arcade.
@@ 1 Introduction
Reinforcement learning (RL) demonstrates remarkable success in solving complex decision-making problems, particularly in fully observable environments (Silver et al., 2016). However, real-world decision-making often involves partial observability, where agents must act based on incomplete or noisy information. Despite its practical importance, partial observability remains understudied in RL research, partially due to the lack of scalable and efficient benchmarks.
Much of the progress in RL relies on standardized benchmarks, such as the Arcade Learning Environment (ALE) (Bellemare et al., 2013) and MuJoCo (Todorov et al., 2012). While these benchmarks have driven significant advancements, they are often fully observable, or become nearly fully observable when considering the last few observations (Machado et al., 2018). This limitation makes it difficult to evaluate algorithms or models that we later apply to realistic, partially observable scenarios. Furthermore, the high sample complexity of modern RL results in the environment bottlenecking the training process (Lu et al., 2024), particularly for on-policy algorithms that require high-throughput simulations (Shacklett et al., 2021).
In this work, we introduce POPGym Arcade, a hardware-accelerated benchmark suite inspired by both POPGym (Morad et al., 2023a) and the ALE. POPGym Arcade contains seven Atari-like environments, with each environment offering both fully observable and partially observable twins. By sharing transition dynamics and observation spaces between all environments and twins, we can isolate and study observability under controlled conditions.
@@
Figure 1: Observations from our environment twins. Observation and actions spaces are shared between all implemented environments, enabling counterfactual studies with observability as the independent variable.
@@ 2 Background
In RL, we interact with an environment to maximize the discounted cumulative reward. A Markov Decision Process (MDP) (Sutton & Barto, 2018) is a tuple $\langle S,A,P,R,\gamma\rangle$, with the set of states $S$, action set $A$, and state transition function $P:S\times A\mapsto\Delta(S)$, where $\Delta(S)$ corresponds to a distribution over the state space. MDPs also contain a reward function $R:S\times A\mapsto\mathbb{R}$ and discount factor $\gamma\in[0,1]$. At each timestep, we receive a state $s_{t}$ and select an action from a policy $a_{t}\sim\pi(s_{t})$ that takes us to the next state following $P(s_{t+1}\mid s_{t},a_{t})$ while emitting a reward $r_{t}=R(s_{t},a_{t})$. We aim to learn the policy $\pi:S\mapsto\Delta(A)$ that maximizes the expected cumulative reward $\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R(s_{t},a_{t})\mid s_{0}=s]$.
@@ 2.1 Partially Observable Markov Decision Processes
In many realistic scenarios, the agent does not have access to the full state of the environment, instead perceiving portions of the state through noisy or limited sensors. We call this problem setup a Partially Observable Markov Decision Process (POMDP). A POMDP is an extension of the MDP, consisting of tuple $\langle S,A,P,R,\gamma,\Omega,O\rangle$, with the set of observations $\Omega$ and the observation function $O:S\mapsto\Delta(\Omega)$. At each timestep, the agent can only access the current observation $o_{t}\sim O(s_{t})$. The observation is not guaranteed to be Markov, thus providing only partial observability of the state, and preventing naive application of standard RL approaches to POMDPs. Fortunately, the sequence of all prior observations and actions, known as the trajectory $\mathbf{x}_{t}=(o_{0},a_{0},o_{1},a_{1}\dots a_{t-1},o_{t})$, is itself Markov (Sutton & Barto, 2018). As such, an optimal policy must consider all prior observations and actions.
@@ 2.2 Memory and Recurrences
It is cumbersome to learn a policy over a variable size $\mathbf{x}_{t}$. We find it simpler to introduce a memory model $f$ that produces a fixed-size latent Markov state $\hat{s}\in\hat{S}$ from the trajectory $\mathbf{x}_{t}$. This confines the complexities associated with partial observability to $f$, allowing our policy to operate on a latent Markov state $\hat{s}$. More formally, a recurrent memory model $f:X\times H\mapsto\hat{S}\times H$ maintains a summary of the trajectory in the recurrent state $h\in H$ and outputs a latent Markov state $\hat{s}\in\hat{S}$ for the policy
@@
Associative recurrent models, also known as linear recurrent models or memoroids (Morad et al., 2024), utilize a recurrent update $\bullet$ that obeys the associative property
$s_{3},h_{3}=(x_{1}\bullet x_{2})\bullet x_{3}=x_{1}\bullet(x_{2}\bullet x_{3}).$ (2)
Given a recurrent associative update, we can leverage the associative property to rearrange the order of operations. For example, given four inputs $x_{1},x_{2},x_{3},x_{4}$, we can compute either
$s_{4},h_{4}=(((x_{1}\bullet x_{2})\bullet x_{3})\bullet x_{4})$ $s_{4},h_{4}=(x_{1}\bullet x_{2})\bullet(x_{3}\bullet x_{4}).$ (3)
The former case corresponds to standard recurrent networks, relying on the prior recurrent state to compute the current state, resulting in linear time complexity. In the latter case, we can compute $x_{1}\bullet x_{2}$ and $x_{3}\bullet$ $x_{4}$ in parallel, achieving logarithmic parallel time complexity (Hinze, 2004). Blelloch (1990) provide an associative scan implementation with linear space complexity.
State-Space Models
$h_{t}=\overline{A}h_{t-1}+\overline{B}x_{t}$ $\hat{s}_{t}=\overline{C}h_{t}+\overline{D}x_{t},$ (4)
where $\overline{A},\overline{B},\overline{C},\overline{D}$ are discretizations of carefully initialized trainable parameters $A,B,C,D$. In practice, we initialize $A,B,C,D$ deterministically. Deterministic initialization applied to many consecutive SSM layers can result in instabilities, and so Orvieto et al. (2023) proposes a meticulously derived random initialization and new parameterization of SSM parameters. They call their method the Linear Recurrent Unit (LRU).
Minimal Recurrent Networks
$h_{t}=(1-\sigma(Ax_{t}+b))\odot h_{t}+\sigma(Ax_{t}+b)\odot\tanh(Cx_{t}+d)$ $\hat{s}_{t}=h_{t},$ (5)
with trainable parameters $A,b,C,d$, sigmoid function $\sigma$, and elementwise product $\odot$. The authors find that the MinGRU outperforms multiple SSM variants across offline reinforcement learning tasks, using a decision transformer (Chen et al., 2021) architecture. We note that the decision transformer optimization objective differs from standard policy gradient or value learning objectives. It is not immediately clear how the MinGRU performs on more traditional value estimation or online RL tasks.
@@ 2.3 JAX and Hardware Acceleration
JAX (Bradbury et al., 2018) is a numerical computing library, supporting automatic differentiation, just-in-time compilation, vectorization, and parallelization. JAX inspects python functions and translates them to XLA bytecode for hardware-specific Just-In-Time (JIT) compilation. XLA compilation targets include CPUs, Tensor Processing Units (TPUs), GPUs and more. The compiler optimizes and fuses XLA operations to reduce latency and improve throughput. In RL, compiling the environment and algorithm together enables optimizations across the environment-agent interface.
@@ 3 Related Work
Progress in RL builds upon a large collection of MDPs and POMDPs (Table 1). Hausknecht & Stone (2015) gauge LSTM on the ALE benchmark, finding that the LSTM underperforms a simpler framestack model that considers the last four observations in almost all environments and suggesting that partial observability is not a big factor in Atari games. Burda et al. (2022) find that even difficult Atari games like Montezuma’s Revenge are close to fully observable. These findings line up with Machado et al. (2018), which introduce difficulties to weaken the observability of Atari. Similarly, Ni et al. (2022) mask positions and velocities in the popular MuJoCo (Todorov et al., 2012) benchmark to induce partial observability. Morad et al. (2023b) find that even memory-free policies can perform well on certain masked control tasks.
We call Atari or masked control tasks that admit near-optimal solutions without memory as weak POMDPs. We also include tasks that require minimal effort to induce full observability (e.g., the last four observations is a Markov state). We are interested in strong POMDPs, where long-term memory is necessary for good performance. More recent work propose strong POMDPs specifically to probe the capabilities of reinforcement learning under partial observability. Such work includes navigation tasks (Chevalier-Boisvert et al., 2018; Beattie et al., 2016; Kempka et al., 2016), psychological experiments (Fortunato et al., 2019), as well as party games (Pleines et al., 2022). Unfortunately, training memory models is expensive, making large scale strong POMDP studies rare (Morad, 2024). Morad et al. (2023a) propose POPGym, a set of POMDPs with low dimensional observation spaces. Unfortunately, the simplified observation spaces make it unclear if POPGym-tested methods will transfer to more complex domains, such as pixel-based tasks. Pixel feature extraction and memory are separate tasks, but combining the two could lead to new emergent capabilities (Reed et al., ).
@@
Copying data from CPU-based environments to GPU-based policies is a major bottleneck in RL (Lu et al., 2022). Using JAX, one can implement, vectorize (duplicate), and compile their environments directly on accelerator hardware. By executing both the training algorithm and environments on-device, users avoid cpu-to-accelerator copies, training their policies up to one thousand times faster (Lange, 2022; Koyamada et al., 2023; Bonnet et al., 2024; Matthews et al., 2024b). We highlight Matthews et al. (2024a); Pignatelli et al. (2024) which offer hardware-accelerated strong POMDPs. Lu et al. (2024) reimplement POPGym in JAX, achieving even greater efficiency.
One may wonder what effect improved sample throughput has on algorithms research. Hessel et al. (2021) propose the Podracer paradigm for policy search, targeting high sample throughput scenarios. Podracer algorithms include variants of policy gradient (Lu et al., 2022) and PPO (Lange, 2022; Lu et al., 2024; Toledo, 2024). In this work, we focus on the PQN algorithm (Gallici et al., 2024), a simplified parallel version of Q learning. Unlike DQN (Mnih et al., 2015), PQN does not use target networks or replay buffers, and opts for on-policy TD($\lambda$) objective.
Table 1: Related benchmarks and their characteristics. A tilde in the MDP Twins columns means that MDP and POMDP twins do not share observation spaces, making causal studies difficult.
@@
MiniGrid (Chevalier-Boisvert et al., 2018)
@@
Navix (Pignatelli et al., 2024)
@@ 4 Approach
In this section, we describe our environment properties and implementation details. All proposed environments have observation space $\Omega=[0,1]^{256\times 256\times 3}$ ($256\times 256$ pixel color images). Similarly, all environments share the action space $A=\{\mathrm{up},\mathrm{down},\mathrm{left},\mathrm{right},\mathrm{fire}\}$. We bound the undiscounted return for all environments in $[-1,1]$ to minimize hyperparameter tuning across tasks. Finally, each environment has three distinct difficulty levels to measure long-term progress in RL.
4.1 Partially and Fully Observable Twins
In our framework, MDP and POMDP environment twins share identical underlying state representations $s\in S$ and transition dynamics $P(s_{t+1}\mid s_{t},a_{t})$, but exhibit distinct observation functions. For an MDP, the observation is guaranteed to be Markov, while this is not the case for a POMDP. The clear boundary between internal state and observation function allows each environment to function as either a MDP or as a POMDP. One can even change observability mid-episode. With our POMDP-MDP twins, we execute studies with observability as the only independent variable, providing stronger evidence and more credibility.
4.2 Vectorized State Transitions and Rendering
We focus on maintaining high environment throughput. First and foremost, we ensure that all environments are both vectorizable and compilable to leverage hardware acceleration. Beyond this, we find a number of small tricks can improve performance. Pre-caching sprites and their positions upon reset provides noticeable performance gains, and replacing calls to jax.lax.cond with jax.lax.switch provides increased throughput. See Appendix F for more information.
@@ 4.3 Memory Introspection Tools
With pixel observations, one can draw conclusions about how pixels impact value predictions (Greydanus et al., 2018). We implement visualization tools to probe which pixels persist in agent memory, and their impact on Q value predictions. More formally, for any timestep we can compute the latent Markov state $\hat{s}$ and recurrent state $h$ from the observation-actions pairs $x$: $\hat{s}_{t},h_{t}=f(x_{t},h_{t-1})$. After training a policy, we initialize $h_{0}$ and scan the greedy max Q policy over the environment $(P,O,R)$, as we would during a rollout: $\hat{s}_{t},\mathbf{x}_{t}=\mathrm{scan}(f,Q,P,O,R)$. Finally, we take the magnitude of the partial derivative of $Q$ values, and backpropagate through time
$\sum_{a_{t}\in A}\left\lvert\frac{\partial}{\partial\mathbf{o}_{t}}Q(\hat{s}_{t},a_{t})\right\rvert\in\Omega^{t}=\sum_{a_{t}\in A}\left\lvert\frac{\partial}{\partial\hat{s}_{t}}Q(\hat{s}_{t},a_{t})\begin{bmatrix}\frac{\partial\hat{s}_{t}}{\partial o_{0}}&\frac{\partial\hat{s}_{t}}{\partial o_{1}}&\dots&\frac{\partial\hat{s}_{t}}{\partial o_{t}}\end{bmatrix}^{\top}\right\rvert.$ (6)
This equation measures how much each pixel from $o_{t-k};k\in\{0\dots t\}$ propagates through memory and contributes to the value at time $t$. Using these tools, we can attribute decisions to specific memories and better understand agent behavior.

### v5

@@ Abstract
We present the POPGym Arcade [^1] Code available at https://github.com/bolt-research/popgym_arcade, a collection of hardware-accelerated, pixel-based environments with shared observation and action spaces. Each environment includes fully and partially observable variants, enabling counterfactual studies on partial observability. We also introduce mathematical tools for analyzing policies under partial observability, which reveal how agents recall past information to make decisions. Our analysis shows (1) that controlling for partial observability is critical and (2) that agents with long-term memory learn brittle policies that struggle to generalize. Finally, we demonstrate that recurrent policies can be “poisoned” by old, out-of-distribution observations, with implications for sim-to-real transfer, imitation learning, and offline reinforcement learning.
@@ 1 Introduction
Real-world decision making often requires reasoning under partial observability, where important information is hidden from the agent. This is a challenge for reinforcement learning (RL), which excels in fully observable settings (Silver et al., 2016). Scalable, efficient benchmarks for partial observability remain scarce, and existing solutions often lack tools to analyze how agents adapt to incomplete information.
Benchmarks like the Arcade Learning Environment (ALE) (Bellemare et al., 2013) and MuJoCo (Todorov et al., 2012) have driven RL progress, but they provide full or near-full observability (Machado et al., 2018), limiting their utility for partially observable scenarios. Prior work also rarely investigates how agents internally compensate for missing information over time.
We introduce POPGym Arcade (Fig. 1), a pixel-based, hardware-accelerated benchmark inspired by POPGym (Morad et al., 2023a) and the ALE. Each environment includes “twins”: fully and partially observable variants with identical dynamics, enabling controlled studies of observability. We complement this with tools to visualize how agents store and recall information, addressing the opacity of decision making under partial observability. By combining hardware acceleration, environment twins, and pixel observations with our visualization tools, we perform efficient counterfactual studies on observability, in a high-dimensional and interpretable observation space.
Contributions
In this work, we contribute:
  • GPU-accelerated, Atari-style partially observable environments.
  • Fully observable twins for counterfactual studies on observability.
  • Tools to examine how agents store and recall information over time.
  • Evidence that solutions to partial observability introduce confounding factors.
  • Evidence that recurrent value functions do not generalize well.
@@
Figure 1: Observations from our environment twins. All environments share a unifying observation, state, and action space, enabling counterfactual studies with observability as the sole independent variable.
@@ 2 Background
In RL, we interact with an environment to maximize the discounted cumulative reward. A Markov Decision Process (MDP) (Sutton & Barto, 2018) is a tuple $\langle S,A,P,R,\gamma\rangle$, with the set of states $S$, actions $A$, and transition function $P:S\times A\mapsto\Delta(S)$, where $\Delta(S)$ is a distribution over $S$. MDPs also contain a reward function $R:S\times A\mapsto\mathbb{R}$ and discount factor $\gamma\in[0,1]$. At each timestep, we receive a state $s_{t}\in S$ and select an action $a_{t}\in A$ from a policy $a_{t}\sim\pi(\cdot|s_{t})$ that takes us to the next state following $P(s_{t+1}\mid s_{t},a_{t})$ while emitting a reward $r_{t}=R(s_{t},a_{t})$. We aim to learn the policy $\pi:S\mapsto\Delta(A)$ that maximizes the expected return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s]$.
@@ 2.1 Partially Observable Markov Decision Processes
In many realistic scenarios, the agent does not have access to the full state of the environment, instead perceiving portions of the state through noisy or limited sensors. We call this problem setup a Partially Observable Markov Decision Process (POMDP). A POMDP is an extension of the MDP, consisting of tuple $\langle S,A,P,R,\gamma,\Omega,O\rangle$, with the set of observations $\Omega$ and the observation function $O:S\mapsto\Delta(\Omega)$. At each timestep, the agent can only access the current observation $o_{t}\sim O(s_{t})$. The observation is not guaranteed to be Markov, thus providing only partial observability of the state, and preventing naive application of standard RL approaches to POMDPs. Fortunately, the sequence of all prior observations and actions, known as the trajectory $\mathbf{x}_{t}=(o_{0},a_{0},o_{1},a_{1}\dots a_{t-1},o_{t})$, is itself Markov (Sutton & Barto, 2018). As such, an optimal policy should consider all prior observations and actions.
@@ 2.2 Memory and Recurrences
It is cumbersome to learn a policy over a variable size $\mathbf{x}_{t}$. We find it simpler to introduce a memory model $f$ that produces a fixed-size latent Markov state $\hat{s}\in\hat{S}$ from the trajectory $\mathbf{x}_{t}$. This confines the complexities associated with partial observability to $f$, allowing the policy to operate on a latent Markov state $\hat{s}$. More formally, a recurrent memory model $f:X\times H\mapsto\hat{S}\times H$ maintains a summary of the trajectory in the recurrent state $h\in H$ and outputs a latent Markov state $\hat{s}\in\hat{S}$ for the policy
@@
Associative recurrent models, also known as linear recurrent models or memoroids (Morad et al., 2024), update the recurrent state with a binary operator $\bullet$ that obeys the associative property
$h_{3}=(x_{1}\bullet x_{2})\bullet x_{3}=x_{1}\bullet(x_{2}\bullet x_{3}).$ (2)
Given an associative recurrent update, we can leverage the associative property to rearrange the order of operations. For example, given four inputs $x_{1},x_{2},x_{3},x_{4}$, we can compute either
$h_{4}=(((x_{1}\bullet x_{2})\bullet x_{3})\bullet x_{4})$ $h_{4}=(x_{1}\bullet x_{2})\bullet(x_{3}\bullet x_{4}).$ (3)
The former case corresponds to standard recurrent networks, relying on the prior recurrent state to compute the current state, resulting in linear time complexity. In the latter case, we can compute $x_{1}\bullet x_{2}$ and $x_{3}\bullet$ $x_{4}$ in parallel, achieving logarithmic parallel time complexity (Hinze, 2004). Blelloch (1990) provide an associative scan implementation with linear space complexity. Associative recurrences are orders of magnitude faster than classical RNNs in practice, while using much less memory than transformers, making them useful in sample inefficient tasks like RL (Lu et al., 2024).
Our experiments rely on four distinct associative recurrent models: the Fast Autoregressive Transformer (FART) (Katharopoulos et al., 2020), a form of State-Space Model (Gu et al., 2022) called the Linear Recurrent Unit (LRU)(Orvieto et al., 2023), and a GRU (Chung et al., 2014) variant of a Minimal Recurrent Network called the MinGRU (Feng et al., 2024). We provide formal descriptions of each model in Appendix H.
@@ 2.3 JAX and Hardware Acceleration
JAX (Bradbury et al., 2018) is a numerical computing library, supporting automatic differentiation, Just-In-Time (JIT) compilation, vectorization, and parallelization. JAX inspects python functions and translates them to XLA bytecode for hardware-specific JIT compilation. XLA compilation targets include CPUs, Tensor Processing Units (TPUs), GPUs and more. In RL, compiling the environment and algorithm together enables optimizations across the environment-agent interface.
@@ 3 Related Work
POMDP Benchmarks
Progress in RL builds upon benchmarks that provide MDPs and POMDPs (Table 1). We specifically differentiate between weak POMDPs and strong POMDPs. Weak POMDPs admit near-optimal agents that consider just a few recent observations. Examples include Atari (Bellemare et al., 2013) which is known to be near-fully observable with framestacking (Machado et al., 2018), even for difficult games like Montezuma’s Revenge (Burda et al., 2022). Velocity-masked MuJoCo (Todorov et al., 2012; Ni et al., 2022) is another such weak POMDP (Morad et al., 2023a). We focus on strong POMDPs, where information is distributed over the trajectory such that long-term memory is necessary to achieve near-optimal returns. Examples of strong POMDPs include navigation tasks (Chevalier-Boisvert et al., 2018; Beattie et al., 2016; Kempka et al., 2016), psychological experiments (Fortunato et al., 2019), and memory games (Pleines et al., 2022; Morad et al., 2023a). Unfortunately, the cost of training memory models combined with the sample inefficiency of RL is prohibitively expensive, forcing researchers into a tradeoff between simplistic low-dimensional observation spaces, and skimping on the number of experiments used to back their claims.
@@
Copying environment data from CPU to GPU during rollouts is a major efficiency bottleneck in the training process (Lu et al., 2022). Using JAX, one can implement their environments directly on accelerator hardware, avoiding cpu-to-accelerator copies and training policies up to one thousand times faster (Lange, 2022; Koyamada et al., 2023; Matthews et al., 2024b). We highlight Matthews et al. (2024a); Pignatelli et al. (2024); Lu et al. (2024) which offer hardware-accelerated strong POMDPs.
With improved environment throughput, we can consider new training paradigms like Podracer (Hessel et al., 2021; Toledo, 2024). In this work, we focus on the PQN algorithm (Gallici et al., 2024), a simplified podracer version of Q learning. Unlike DQN (Mnih et al., 2015), PQN does not use target networks or replay buffers, and opts for an on-policy TD($\lambda$) objective.
Analyzing Memory in RL
Prior work often focuses on designing memory models that maximize the return, but disentangling memory efficacy from confounding factors requires controlled benchmarks. Performance gaps between memory models could stem from parameter count disparities (Hansen et al., 2023) or implicit inductive biases (Morad et al., 2023b), rather than the ability of a memory model to learn a latent Markov state. To isolate the impact of partial observability, fully observable experimental controls (MDP twins) are critical. They allow us to quantify how well a memory model mitigates partial observability with certainty. Without such controls, claims about memory performance remain speculative. Controls that span environments are rare in literature – masked MuJoCo (Ni et al., 2022), for example, does not use a unified observation and action space between environments, necessitating changes in model architecture and weakening the control. To date, we are only aware of two benchmarks that enable controlling for observability across environments (Pleines et al., 2022; Rajan et al., 2023), and neither utilize hardware acceleration.
By measuring the return, we indirectly measure the quality of the latent Markov state $\hat{s}$ created by the memory model. However, few works directly answer how the model constructs this latent state, providing hints but never complete answers. Hausknecht & Stone (2015); Kapturowski et al. (2019) measure how memory models function with stale recurrent states, while Morad et al. (2024) suggest that recurrent value functions may not operate as intended. Ni et al. (2024) decouples memory performance from temporal credit assignment, more directly measuring the quality of $\hat{s}$. Finally, (Elelimy et al., 2024; Morad et al., 2023b) examine the distribution of recurrent states. There is a gap in understanding concerning the meaning or source of memory contents.
Table 1: Related benchmarks and their characteristics. A tilde in the MDP Twins columns means that MDP and POMDP twins do not share observation spaces, making causal studies difficult. Our benchmark is the only one with both strong POMDP/MDP twins and GPU acceleration.
@@
MiniGrid (Chevalier-Boisvert et al., 2018) $\sim$
@@
Navix (Pignatelli et al., 2024) $\sim$
Brax/PixelBrax
(Freeman et al., 2021; McInroe & Garcin, 2025)
@@ 4 Approach
In this section, we propose POPGym Arcade and derive tools for memory visualization.
4.1 Environments
POPGym Arcade consists of seven base environments (Appendix E), with a total of 84 environment configurations. Each environment has both small and large pixel observation spaces $\Omega_{\mathrm{small}}=[0,1]^{128\times 128\times 3},\Omega_{\mathrm{large}}=[0,1]^{256\times 256\times 3}$. Similarly, all environments share the action space $A=\{\mathrm{up},\mathrm{down},\mathrm{left},\mathrm{right},\mathrm{fire}\}$. We bound the undiscounted return for all environments in $[-1,1]$ to simplify optimality gap and related comparisons. Finally, we implement three distinct difficulties for each environment.
Partially and Fully Observable Twins
In our framework, MDP and POMDP environment twins share identical underlying state representations $s\in S$ transition dynamics $P(s_{t+1}\mid s_{t},a_{t})$, and observation spaces $o\in\Omega$, but exhibit distinct observation functions $O$. For the MDP variants, every observation $o$ is guaranteed to be Markov, which is not true for the POMDP variants. We enforce an API boundary between $P$ and $O$, which allows each environment to function as either a MDP or as a POMDP. A single shared observation and action space means we can use identical agent configurations for both POMDPs and MDPs. One can even change the observability or game itself mid-episode. With our POMDP-MDP twins, we execute studies with observability as the only independent variable. Model architecture, hyperparameters, and everything else can remain fixed.
Vectorized State Transitions and Rendering
Given the poor sample efficiency of RL, we focus on maintaining high environment throughput. First and foremost, we ensure that all environments are both vectorizable and compilable to leverage hardware acceleration. Beyond this, we find a number of small tricks can improve performance. Pre-caching sprites and their pixel positions upon reset provides noticeable performance gains, and replacing calls to jax.lax.cond with jax.lax.switch provides increased throughput. See Appendix F for more information.
@@ 4.2 Memory Introspection Tools
With pixel observations, one can draw conclusions about how pixels impact value predictions (Greydanus et al., 2018). We implement visualization tools to probe which pixels the agent remembers, and their impact on the policy. More formally, given a trajectory $\textbf{x}_{n}$, we compute latent Markov states $\hat{s}_{0},\hat{s}_{1},\dots,\hat{s}_{n}$ via Eq. 1. Then, we backpropagate through memory and policy, taking the absolute value of the gradient of $Q$ values with respect to an observation $o_{t}$ where $t\leq n$
$\sum_{a_{n}\in A}\left\lvert\nabla_{o_{t}}Q(\hat{s}_{n},a_{n})\right\rvert=\sum_{a_{n}\in A}\left\lvert\left(\frac{\partial}{\partial\hat{s}_{n}}Q(\hat{s}_{n},a_{n})\right)\frac{\partial\hat{s}_{n}}{\partial o_{t}}\right\rvert,$ (4)
where ||$\cdot$|| ⋅ | denotes elementwise absolute value. This equation measures how much each pixel from a prior observation $o_{t}$ propagates through memory and contributes to the $Q$ values at time $n$. We derive a similar equation for policy gradient methods, instead measuring the impact on the action distribution
[…]
These methods produce qualitative saliency maps, describing pixel-level memory recall. We are also interested in a more quantitative and consistent measure, that elucidates the inner workings of memory in expectation over a distribution of trajectories. For these cases, we introduce the recall density, which we derive below.
Recall Density
To achieve a more general notion of memory saliency, we start with Eqs. 4 and 5 and reduce over the observation dimensions using the $L_{1}$ norm, such that we receive a scalar time-varying function $z$
[…]
where $z_{Q}$ represents the quantity for $Q$ functions and $z_{\pi}$ for policies. Then, we normalize $z$, representing a scale-invariant density over trajectory $\mathbf{x}_{n}$
$\delta_{Q}(\mathbf{x}_{n},t)=\frac{z_{Q}(\mathbf{x}_{n},t)}{\sum_{i=0}^{n}z_{Q}(\mathbf{x}_{n},i)}$ $\delta_{\pi}(\mathbf{x}_{n},t)=\frac{z_{\pi}(\mathbf{x}_{n},t)}{\sum_{i=0}^{n}z_{\pi}(\mathbf{x}_{n},i)}.$ (7)
Finally, we take the sample mean over $m$ trajectories to approximate the expected density under a policy $\pi$ and memory model $f$
$\frac{1}{m}\sum_{i=1}^{m}\delta_{Q}(\mathbf{x}^{(i)}_{n},t)\approx\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x}_{n},t)]$ $\frac{1}{m}\sum_{i=1}^{m}\delta_{\pi}(\mathbf{x}^{(i)}_{n},t)\approx\mathbb{E}_{\pi,f}[\delta_{\pi}(\mathbf{x}_{n},t)].$ (8)
We may compute the resulting expectation over all $0\leq t\leq n$ to generate a plot summarizing memory recall.
Generalized Measures
Thus far, $n$ must be equal across all sampled trajectories, while in practice trajectories often differ in length. We use generalized time coordinates to resolve this issue. For each trajectory $\mathbf{x}^{(i)}$ of length $n_{i}$, we normalize the timestep $t\in\{0,\dots,n_{i}\}$ to a value $\tau=t/n_{i}\in[0,1]$. We then compute the sample mean of the empirical densities at corresponding normalized time points across $m$ trajectories
$\frac{1}{m}\sum_{i=1}^{m}\delta_{Q}(\mathbf{x}^{(i)}_{n_{i}},\lfloor\tau\cdot n_{i}\rfloor)\approx\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x},\tau)]$ $\frac{1}{m}\sum_{i=1}^{m}\delta_{\pi}(\mathbf{x}^{(i)}_{n_{i}},\lfloor\tau\cdot n_{i}\rfloor)\approx\mathbb{E}_{\pi,f}[\delta_{\pi}(\mathbf{x},\tau)].$ (9)
We may compute the expectation over all $\tau\in[0,1]$ to generate a plot summarizing memory recall. One should ensure that $\mathbf{x}$ is not terminal, or risk evaluating $Q$ or $\pi$ at a terminal state.
Definition 4.1 (Recall Density).
The Recall Density measures the expected relative influence of a past observation at the current timestep. For trajectories generated by policy $\pi$ and memory model $f$, it quantifies how much an observation at normalized time $\tau\in[0,1]$ impacts either:
  • The $Q$ value at trajectory end ($\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x},\tau)]$), via $Q$ gradient magnitudes
  • The action distribution at trajectory end ($\mathbb{E}_{\pi,f}[\delta_{\pi}(\mathbf{x},\tau)]$), via policy gradient magnitudes
Both variants normalize each observation’s influence by its trajectory’s total gradient magnitude, then average across trajectories. This handles variable-length trajectories by aligning timesteps to $\tau$.